Juanita goes hiking every weekend,while Steve usually stays in the city to drink coffee and read the newspaper.Both Juanita and Steve decide to climb Mt. Boulderdash on the same weekend in August,but do not hike it together.Juanita can hike 3 times as fast as Steve and the hike takes them a combined total of 9 hours.How long does it take Juanita to climb the mountian?
Answer by checkley71(8403)

X+3X=9
4X=9
X=9/4
X=2.25 HOURS IS JUANITA'S CLIMB TIME
PROOF
2.25+3*2.25=9
2.25+6.75=9
9=9